A company needs to have $105,000 in 5 years, and will create a fund to insure that the $105,000 will be available. If it can earn a 6% return compounded annually, how much must the company invest in the fund today to equal the $105,000 at the end of 5 years? (PV of $1, FV of $1, PVA of $1, and FVA of $1) (Use appropriate factor(s) from the tables provided.)

Multiple Choice

  • $78,467

  • $98,700

  • $140,506

  • $31,500

  • $73,500
Answer #1

Calculate initial investment

Future value = Present value*(1+i)n

105000 = X*(1+.06)5

X = 105000/1.33823 = 78467

So answer is a) $78467